OpenStudy (anonymous):

A,b,c shared 92 cookies. A got 2 cookies more than B. When B got was 4 times as much as C's. How many cookies did C get?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

how can i solve a+b+c?

OpenStudy (anonymous):

yyyyyyyyyyyyyyesssssss

OpenStudy (anonymous):

yeesssssss baaaby

OpenStudy (anonymous):

ok help me D:

OpenStudy (anonymous):

nononnonononono i don't know whaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaat to do

OpenStudy (anonymous):

likkkkkkke a somebooooody

OpenStudy (anonymous):

caaaaaan sooomebooody teeeach me plleeeeeeeeeeeeasse?

OpenStudy (anonymous):

Let A, b and c be the number of cookies each received respectively. Solve the following for A, b and c.\[\{A=b+2,b=4 c,A+b+c=92\} \]\[\{A=42,b=40,c=10\} \]